# Large Language Model for Vulnerability Detection and Repair: Literature Review and Roadmap

### v1

@@ Abstract.
The significant advancements in Large Language Models (LLMs) have resulted in their widespread adoption across various tasks within Software Engineering (SE), including vulnerability detection and repair. Numerous recent studies have investigated the application of LLMs to enhance vulnerability detection and repair tasks. Despite the increasing research interest, there is currently no existing survey that focuses on the utilization of LLMs for vulnerability detection and repair. In this paper, we aim to bridge this gap by offering a systematic literature review of approaches aimed at improving vulnerability detection and repair through the utilization of LLMs. The review encompasses research work from leading SE, AI, and Security conferences and journals, covering 36 papers published at 21 distinct venues. By answering three key research questions, we aim to (1) summarize the LLMs employed in the relevant literature, (2) categorize various LLM adaptation techniques in vulnerability detection, and (3) classify various LLM adaptation techniques in vulnerability repair. Based on our findings, we have identified a series of challenges that still need to be tackled considering existing studies. Additionally, we have outlined a roadmap highlighting potential opportunities that we believe are pertinent and crucial for future research endeavors.
copyright: none
@@ 1. Introduction
A software vulnerability refers to a flaw or weakness in a software system that can be exploited by attackers. Recently, the number of software vulnerabilities has increased significantly (TARGETT, 2022), affecting numerous software systems. To mitigate these issues, researchers have proposed methods for automatic detection and repair of identified vulnerabilities. However, traditional techniques, such as rule-based detectors or program analysis-based repair tools, encounter challenges due to high false positive rates (Shahriar and Zulkernine, 2012) and their inability to work for diverse types of vulnerabilities (Zhou et al., 2024a), respectively.
@@
Recently, Large Language Models (LLMs) pre-trained on large corpus have demonstrated remarkable effectiveness across various natural language and software engineering tasks (Hou et al., 2023). Given their recent success, researchers have proposed various LLM-based approaches to improve automated vulnerability detection and repair, demonstrating promising outcomes for both detection and repair tasks (Steenhoek et al., 2023; Zhou et al., 2024a). LLM-based approaches for vulnerability detection and repair are increasingly attracting attention due to their potential to automatically learn features from known vulnerabilities and find/fix unseen ones. Furthermore, LLMs have the potential to utilize rich knowledge acquired from large-scale pre-training to enhance vulnerability detection and repair.
@@
Despite the increasing research interest in utilizing LLMs for vulnerability detection and repair, to the best of our knowledge, there has been no comprehensive literature review summarizing on the state-of-the-art approaches, current challenges, and future directions in this field. To effectively chart the most promising path forward for research on the utilization of LLM techniques in vulnerability detection and repair, we conducted a Systematic Literature Review (SLR) to bridge this gap, providing valuable insights to the community. In this paper, we collected 36 primary studies over the last 6 years (2018-2024). We then summarized the LLMs used by relevant papers and classified various techniques used to adapt LLMs. We also discussed the key challenges associated with using LLMs for vulnerability detection and repair and proposed multiple potential research directions.
@@
Related Literature Reviews. Researchers have undertaken a series of research endeavors concerning Machine Learning (ML) for source code vulnerability detection or repair (Hou et al., 2023; Ghaffarian and Shahriari, 2017; Lin et al., 2020; Wu et al., 2022; Zhang et al., 2023a). Hou et al. (Hou et al., 2023) conducted a systematic literature review on LLMs for SE. However, due to the extensive range of SE tasks to summarize, their review did not categorize the detailed utilization of LLMs in vulnerability detection and repair. In contrast, our literature review is more focused on vulnerability detection and repair. Ghaffarian et al. (Ghaffarian and Shahriari, 2017) studied vulnerability analysis and discovery approaches published till 2016. Lin et al. (Lin et al., 2020) reviewed vulnerability detection approaches utilizing deep learning till 2020. Wu et al. (Wu et al., 2022) investigated vulnerability detection approaches published before May 2022. However, their focus was not on LLMs and primarily covered general ML models such as recurrent neural networks. In contrast, our literature review includes papers published until March 2024 and focuses on LLMs, encompassing 15 distinct LLMs utilized in 36 relevant studies. Recently, Zhang et al. (Zhang et al., 2023a) reviewed learning-based automated program repair, which covered vulnerability repair as part of its study scope. In Different from Zhang et al., our literature review is more focused on LLMs and vulnerability repair: we included more recent studies, offered a detailed categorization of LLM usages in vulnerability repair, and discussed specialized future directions for vulnerability repair with LLMs.
@@
  • We present a systematic review of recent 36 primary studies focusing on the utilization of LLMs for vulnerability detection and repair.
  • We offer a comprehensive summary of the LLMs utilized in the relevant literature and categorize various techniques used to adapt LLMs to do these two tasks (vulnerability detection and repair).
  • We discuss the key challenges associated with using LLMs for vulnerability detection and repair, and propose multiple potential research directions.
2. Preliminaries
@@ 2.2. Large Language Models (LLMs)
The term Large Language Model (LLM) was introduced to distinguish language models based on their parameter size, specifically referring to large-sized pre-trained language models (Zhao et al., 2023). However, the literature lacks a formal consensus on the minimum parameter scale for LLMs (Wang et al., 2023). In this paper, we adopt the LLM scope division and taxonomy introduced by Pan et al. (Pan et al., 2023) and categorize the mainstream LLMs into three groups according to their architectures: 1) encoder-only, 2) encoder-decoder, and 3) decoder-only LLMs. We will provide a brief introduction to some representative LLMs for each category due to the space limit.
Encoder-only LLMs. Encoder-only LLMs are a type of neural network architecture that utilizes only the encoder component of the Transformer model (Devlin et al., 2018). In the SE domain, examples of encoder-only LLMs include CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2020), CuBERT (Kanade et al., 2019), VulBERTa (Hanif and Maffeis, 2022), CCBERT (Zhou et al., 2023b), SOBERT (He et al., 2023), and BERTOverflow (Tabassum et al., 2020).
Encoder-decoder LLMs. Encoder-decoder LLMs integrate both the encoder and decoder modules of the Transformer model (Vaswani et al., 2017). The encoder processes the input sentence, while the decoder generates the target output text/code. Prominent examples of encoder-decoder LLMs include PLBART (Ahmad et al., 2021), T5 (Raffel et al., 2020), CodeT5 (Wang et al., 2021), UniXcoder (Guo et al., 2022), and NatGen (Chakraborty et al., 2022).
Decoder-only LLMs. Decoder-only LLMs exclusively utilize the decoder module of the Transformer model to generate the target output text/code. The GPT series, including GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), GPT-3.5 (OpenAI, 2022), and GPT-4 (OpenAI, 2023), stand as prominent implementations of this model series. Additionally, in the SE domain, there are numerous decoder-only LLMs specialized for code as well. Examples include CodeGPT (Lu et al., 2021), Codex (Chen et al., 2021), Polycoder (Xu et al., 2022), Incoder (Fried et al., 2023), CodeGen series (Nijkamp et al., 2023c, a), Copilot (GitHub, 2023), Code Llama (Meta, 2023), and StarCoder (Li et al., 2023a).
@@ 3.1. Research Question
In this paper, we focus on investigating the following three Research Questions (RQs):
@@ 3.2. Search Strategy
As shown in Fig. 1, following the guide by Zhang et al. (Zhang et al., 2011), our initial step is to identify primary studies to answer the research questions (RQs) above. Because the first LLM (i.e., BERT (Devlin et al., 2018)) was introduced in 2018, our search focused on papers published from 2018 onwards (i.e., from January 1st, 2018, to March 1st, 2024). Next, we identified the top peer-reviewed and influential conference and journal venues in the domains of SE, AI, and Security. We included 13 conferences (ICSE, ESEC/FSE, ASE, ISSTA, CCS, S&P, USENIX Security, NDSS, AAAI, IJCAI, ICML, NIPS, ICLR) and 4 journals (TOSEM, TSE, TDSC, TIFS). After the manual searching, we identified 35 papers that were relevant to our research objectives.
In addition to manually searching primary studies from top-tier venues, we also conducted an automated search across 7 popular databases, including IEEE Xplore (IEE, [n. d.]), ACM Digital Library (ACM, [n. d.]), SpringerLink (Spr, [n. d.]), Wiely (Wie, [n. d.]), ScienceDirect (Sci, [n. d.]), Web of Science (Web, [n. d.]) and arXiv (arX, [n. d.]). The search string used in the automated search is crafted from the relevant papers identified in the manual search. Please kindly check the complete set of search keywords in our online appendix (app, 2024) due to the limited space. After conducting the automatic search, we collected 1,500 relevant studies with the automatic search from these 7 popular databases.
@@ 3.3. Study Selection
  • The paper must be a peer-reviewed full research paper published either in a conference proceeding or a journal.
@@
In the first phase, by filtering out short papers (exclusion criteria 1) and deduplication (exclusion criteria 4), the total number of included papers was reduced to 1,332. In the second phase, we manually examined the venue, title, and abstracts of the papers, and the total number of included papers declined to 212. Books, keynote records, panel summaries, technical reports, theses, tool demo papers, editorials, literature reviews, or survey papers were also discarded in this phase (exclusion criteria 2-3). In the third phase, we manually read the full text of the paper to remove irrelevant papers. Specifically, the vulnerability detection or repair papers, which do not utilize LLMs but other methods, e.g., graph neural networks (GNN) or recurrent neural networks (RNN), were dropped (exclusion criteria 5). We also excluded studies that do not focus on source code vulnerability, such as those on binary code, protocols, or network communication. Furthermore, we removed studies that just discussed LLM as an idea or future work (exclusion criteria 6). We also removed the papers focusing on other tasks rather than vulnerability detection or repair, such as vulnerable data generation, vulnerability assessment, etc. (exclusion criteria 7). After the third phase, we identify 61 primary studies directly relevant to our research topic.
@@
Quality Assessment. To prevent biases introduced by low-quality studies, we formulated five Quality Assessment Criteria (see our online appendix (app, 2024)) to assess the clarity, validity, and significance of the papers. We employed a scoring system ranging from 0 to 3 (poor, fair, good, excellent) for each quality assessment criterion. Following the manual assignment of scores, we selected papers with total scores exceeding 12 (80% of the maximum possible score). After this quality assessment, we obtained 33 papers.
Forward and Backward Snowballing. To avoid omitting any possibly relevant work during our manual and automated search process, we also performed lightweight backward and forward snowballing  (Wohlin, 2014). This involved reviewing both the references cited in our selected 33 primary studies and the publications that cited these studies. As a supplement, we gathered 77 more papers and repeated the entire study selection process, including filtering, deduplication, and quality assessment, which resulted in the identification of three additional papers. Thus, we obtained a final set of 36 papers to study.
@@ 3.4. Data Extraction and Analysis
Fig. 2 illustrates the distribution of selected primary studies across each year. The earliest relevant study we identified was published in 2021 (Ziems and Wu, 2021). Subsequently, interest in exploring LLMs for vulnerability detection and repair has steadily increased, peaking in 2023, with 50% of the total studied papers. Additionally, it is noteworthy that in the first two months of 2024 alone, the number of published papers has already reached 10, constituting 27.8% of the total papers. This suggests a growing research interest in leveraging LLMs for vulnerability detection and repair. We also analyzed the publication venues of the included studies. ICSE emerges as the predominant conference venue for LLM studies on vulnerability detection and repair, contributing 27.8% of the total number of studies. Other notable venues are TSE (8.3%), FSE (5.6%), EMSE (5.6%), ISSTA (5.6%), SCAM (5.6%), and TOSEM (2.8%).

### v3

# Large Language Model for Vulnerability Detection and Repair: Literature Review and the Road Ahead

@@ Abstract.
The significant advancements in Large Language Models (LLMs) have resulted in their widespread adoption across various tasks within Software Engineering (SE), including vulnerability detection and repair. Numerous studies have investigated the application of LLMs to enhance vulnerability detection and repair tasks. Despite the increasing research interest, there is currently no existing survey that focuses on the utilization of LLMs for vulnerability detection and repair. In this paper, we aim to bridge this gap by offering a systematic literature review of approaches aimed at improving vulnerability detection and repair through the utilization of LLMs. The review encompasses research work from leading SE, AI, and Security conferences and journals, encompassing 43 papers published across 25 distinct venues, along with 15 high-quality preprint papers, bringing the total to 58 papers. By answering three key research questions, we aim to (1) summarize the LLMs employed in the relevant literature, (2) categorize various LLM adaptation techniques in vulnerability detection, and (3) classify various LLM adaptation techniques in vulnerability repair. Based on our findings, we have identified a series of limitations of existing studies. Additionally, we have outlined a roadmap highlighting potential opportunities that we believe are pertinent and crucial for future research endeavors.
Literature review, vulnerability detection, vulnerability repair, large language models
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Security and privacy Software and application security
@@ 1. Introduction
A software vulnerability refers to a flaw or weakness in a software system that can be exploited by attackers. Recently, the number of software vulnerabilities has increased significantly (security_report2, ), affecting numerous software systems. To mitigate these issues, researchers have proposed methods for automatic detection and repair of identified vulnerabilities. However, traditional techniques, such as rule-based detectors or program analysis-based repair tools, encounter challenges due to high false positive rates (shahriar2012mitigating, ) and their inability to work for diverse types of vulnerabilities (zhou2024out, ), respectively.
@@
Recently, Large Language Models (LLMs) pre-trained on large corpus have demonstrated remarkable effectiveness across various natural language and software engineering tasks (Hou2023LargeLM, ). Given their recent success, researchers have proposed various LLM-based approaches to improve automated vulnerability detection and repair, demonstrating promising outcomes for both detection and repair tasks (steenhoek2023empirical, ; zhou2024out, ). LLM-based approaches for vulnerability detection and repair are increasingly attracting attention due to their potential to automatically learn features from known vulnerabilities and find/fix unseen ones. Furthermore, LLMs have the potential to utilize rich knowledge acquired from large-scale pre-training to enhance vulnerability detection and repair.
@@
Despite the increasing research interest in utilizing LLMs for vulnerability detection and repair, to the best of our knowledge, no comprehensive literature review has yet summarized the state-of-the-art approaches, identified the limitations of current studies, and proposed future research directions in this field. To effectively chart the most promising path forward for research on the utilization of LLM techniques in vulnerability detection and repair, we conducted a Systematic Literature Review (SLR) to bridge this gap, providing valuable insights to the community. In this paper, we collected 58 primary studies over the last 6 years (2018-2024). We then summarized the LLMs used in these studies, classified various techniques for adapting LLMs, and discussed the limitations of existing research. Finally, we proposed a roadmap outlining future research directions and opportunities in this area.
[…]
Figure 1. Structure of This Survey
@@
Related Literature Reviews. Researchers have undertaken a series of research endeavors concerning Machine Learning (ML) for source code vulnerability detection or repair (Hou2023LargeLM, ; ghaffarian2017software, ; lin2020software, ; wu2022code, ; zhang2023surveyrepair, ). Hou et al. (Hou2023LargeLM, ) conducted a systematic literature review on LLMs for SE. Similarly, Zhang et al. (zhang2023unifying, ) systematically reviewed the recent advancements in LLMs for SE. However, due to the extensive range of SE tasks to summarize, their review did not categorize the detailed utilization of LLMs in vulnerability detection and repair. In contrast, our literature review is more focused on vulnerability detection and repair. Ghaffarian et al. (ghaffarian2017software, ) studied vulnerability analysis and discovery approaches published till 2016. Lin et al. (lin2020software, ) reviewed vulnerability detection approaches utilizing deep learning till 2020. Wu et al. (wu2022code, ) investigated vulnerability detection approaches published before May 2022. However, their focus was not on LLMs and primarily covered general ML models such as recurrent neural networks. In contrast, our literature review includes papers published until March 2024 and focuses on LLMs, encompassing 37 distinct LLMs utilized in 58 relevant studies. Recently, Zhang et al. (zhang2023survey, ) reviewed learning-based automated program repair, which covered vulnerability repair as part of its study scope. In Different from Zhang et al., our literature review is more focused on LLMs and vulnerability repair: we included more recent studies, offered a detailed categorization of LLM usages in vulnerability repair, and discussed specialized future directions for vulnerability repair with LLMs.
@@
  • We present a systematic review of recent 58 primary studies focusing on the utilization of LLMs for vulnerability detection and repair.
  • We offer a comprehensive summary of the LLMs utilized in the relevant literature and categorize various techniques used to adapt LLMs to do the two tasks (vulnerability detection and repair).
  • We discuss the limitations of existing studies on using LLMs for vulnerability detection and repair and propose a roadmap outlining future research directions and opportunities.
Survey Structure. Figure 1 outlines the structure of this paper. Section 2 provides the background information, while Section 3 details the methodology. Section 4 discusses the LLMs employed in vulnerability detection and repair. Sections 5 and 6 cover the relevant work on adapting and enhancing LLMs for vulnerability detection and repair, respectively. Section 7 discusses the characteristics of datasets and deployment in LLM-based vulnerability detection and repair studies. Finally, Section 8 discusses the limitations of current studies and presents a roadmap for future research and opportunities.
2. Background and Preliminaries
@@ 2.2. Large Language Models (LLMs)
The term Large Language Model (LLM) was introduced to distinguish language models based on their parameter size, specifically referring to large-sized pre-trained language models (zhao2023survey, ). However, the literature lacks a formal consensus on the minimum parameter scale for LLMs (wang2023software, ). In this paper, we adopt the LLM scope division and taxonomy introduced by Pan et al. (pan2023unifying, ) and categorize the mainstream LLMs into three groups according to their architectures: 1) encoder-only, 2) encoder-decoder, and 3) decoder-only LLMs. We will provide a brief introduction to some representative LLMs for each category due to the space limit.
Encoder-only LLMs. Encoder-only LLMs are a type of neural network architecture that utilizes only the encoder component of the Transformer model (devlin2018bert, ). In the SE domain, examples of encoder-only LLMs include CodeBERT (feng2020codebert, ), GraphCodeBERT (guo2020graphcodebert, ), CuBERT (Kanade2019LearningAE, ), VulBERTa (DBLP:conf/ijcnn/HanifM22, ), CCBERT (zhou2023ccbert, ), SOBERT (he2023representation, ), and BERTOverflow (tabassum2020code, ).
Encoder-decoder LLMs. Encoder-decoder LLMs integrate both the encoder and decoder modules of the Transformer model (vaswani2017attention, ). The encoder processes the input sentence, while the decoder generates the target output text/code. Prominent examples of encoder-decoder LLMs include PLBART (ahmad2021unified, ), T5 (raffel2020exploring, ), CodeT5 (wang2021codet5, ), UniXcoder (DBLP:conf/acl/GuoLDW0022, ), and NatGen (DBLP:conf/sigsoft/ChakrabortyADDR22, ).
Decoder-only LLMs. Decoder-only LLMs exclusively utilize the decoder module of the Transformer model to generate the target output text/code. The GPT series, including GPT-2 (radford2019language, ), GPT-3 (brown2020language, ), GPT-3.5 (openai2022gpt, ), and GPT-4 (openai2023gpt4, ), stand as prominent implementations of this model series. Additionally, in the SE domain, there are numerous decoder-only LLMs specialized for code as well. Examples include CodeGPT (lu2021codexglue, ), Codex (chen2021evaluating, ), Polycoder (DBLP:conf/pldi/Xu0NH22, ), Incoder (DBLP:conf/iclr/FriedAL0WSZYZL23, ), CodeGen series (DBLP:conf/iclr/NijkampPHTWZSX23, ; DBLP:journals/corr/abs-2305-02309, ), Copilot (github2021copilot, ), Code Llama (Meta2023Codellama, ), and StarCoder (Li2023StarCoderMT, ).
@@ 3.1. Research Question
In this paper, we focus on investigating four Research Questions (RQs). The first three RQs mainly focus on identifying which LLMs are used and how they are adapted for vulnerability detection and repair:
@@
The first three RQs pertain primarily to the model construction phase. Generally, the pipeline of a learning-based method can be divided into three phases: 1) data preparation, 2) model construction, and 3) deployment. We also investigate the characteristics of the datasets and deployment designs in LLM-based vulnerability detection and repair studies:
  • RQ4: What are the characteristics of the datasets and deployment strategies in LLM-based vulnerability detection and repair studies?
@@ 3.2. Search Strategy
As shown in Fig. 2, following the guide by Zhang et al. (DBLP:journals/infsof/ZhangBT11, ), our initial step is to identify primary studies to answer the research questions (RQs) above. Because the first LLM (i.e., BERT (devlin2018bert, )) was introduced in 2018, our search focused on papers published from 2018 onwards (i.e., from January 2018, to March 2024). Next, we identified the top peer-reviewed and influential conference and journal venues in the domains of SE, AI, and Security. We included 13 conferences (ICSE, ESEC/FSE, ASE, ISSTA, CCS, S&P, USENIX Security, NDSS, AAAI, IJCAI, ICML, NIPS, ICLR) and 4 journals (TOSEM, TSE, TDSC, TIFS). After the manual searching, we identified 35 papers that were relevant to our research objectives.
In addition to manually searching primary studies from top-tier venues, we also conducted an automated search across 7 popular databases, including IEEE Xplore (IEEEXplore, ), ACM Digital Library (ACMDigital, ), SpringerLink (SpringerLink, ), Wiely (Wiely, ), ScienceDirect (ScienceDirect, ), Web of Science (WebofScience, ) and arXiv (arXiv, ). The search string used in the automated search is crafted from the relevant papers identified in the manual search. Please kindly check the complete set of search keywords in our online appendix (appendix, ) due to the limited space. After conducting the automatic search, we collected 1,500 relevant studies with the automatic search from these 7 popular databases.
@@ 3.3. Study Selection
  • The paper must be a peer-reviewed full research paper published either in a conference proceeding or a journal.  [^1] For preprint papers released on arXiv in 2023 or 2024 that are not yet published due to their recent release, we will retain them when checking the venues and will perform a manual quality check to decide whether to include them in our study.
@@
In the first phase, by filtering out short papers (exclusion criteria 1) and deduplication (exclusion criteria 4), the total number of included papers was reduced to 1,332. In the second phase, we manually examined the venue, title, and abstracts of the papers, and the total number of included papers declined to 212. Please kindly note that, in this step, we retain preprint papers released on arXiv in 2023 or 2024 that are not yet published due to their recent release. Books, keynote records, panel summaries, technical reports, theses, tool demo papers, editorials, literature reviews, or survey papers were also discarded in this phase (exclusion criteria 2-3). In the third phase, we manually read the full text of the paper to remove irrelevant papers. Specifically, the vulnerability detection or repair papers, which do not utilize LLMs but other methods, e.g., graph neural networks (GNN) or recurrent neural networks (RNN), were dropped (exclusion criteria 5). We also excluded studies that do not focus on source code vulnerability, such as those on binary code, protocols, or network communication. Furthermore, we removed studies that just discussed LLM as an idea or future work (exclusion criteria 6). We also removed the papers focusing on other tasks rather than vulnerability detection or repair, such as vulnerable data generation, vulnerability assessment, etc. (exclusion criteria 7). After the third phase, we identify 82 primary studies directly relevant to our research topic.
@@
Quality Assessment. To prevent biases introduced by low-quality studies, we formulated five Quality Assessment Criteria (QAC):
  • QAC1: Was the study published in a prestigious venue?
  • QAC2: Does the study make a contribution to the academic or industrial community?
  • QAC3: Does the study provide a clear description of the workflow and implementation of the proposed approach?
  • QAC4: Are the experiment details, including datasets, baselines, and evaluation metrics, clearly outlined?
  • QAC5: Do the findings from the experiments strongly support the main arguments presented in the study?
We employed a scoring system ranging from 0 to 3 (poor, fair, good, excellent) for each quality assessment criterion. Following the manual assignment of scores, we selected papers with total scores reaching 12 (80% of the maximum possible score). Please note that for preprint papers released on arXiv in 2023 or 2024 that are not yet published, their score for the “venue” criterion is 0. However, if their total score reaches 12, we still consider them in our study. After this quality assessment, we obtained 42 papers. Among them, 37 papers were published, and 5 high-quality preprint papers were included after passing the quality assessment.
Forward and Backward Snowballing. To avoid omitting any possibly relevant work during our manual and automated search process, we also performed lightweight backward and forward snowballing  (DBLP:conf/ease/Wohlin14, ). This involved reviewing both the references cited in our selected 42 primary studies and the publications that cited these studies. As a supplement, we gathered 83 more papers and repeated the entire study selection process, including filtering, deduplication, and quality assessment, which resulted in the identification of 16 additional papers. Thus, we obtained a final set of 58 papers to study.
@@ 3.4. Data Extraction and Analysis
Fig. 3 illustrates the distribution of selected primary studies across each year. The earliest relevant study we identified was published in 2021 (DBLP:conf/infocom/ZiemsW21, ). Subsequently, interest in exploring LLMs for vulnerability detection and repair has steadily increased, peaking in 2024, with 46.6% of the total studied papers. The growing number of papers on these topics indicates increasing research interest in leveraging LLMs for vulnerability detection and repair. We also analyzed the publication venues of the included studies. ICSE emerges as the predominant conference venue for LLM studies on vulnerability detection and repair, contributing 20.7% of the total number of studies. Other notable venues are TSE (5.2%), FSE (3.4%), EMSE (3.4%), ISSTA (3.4%), MSR (3.4%), and TOSEM (1.7%).